OpenStudy (anonymous):

What is the height of a cone with a radius of 2 inches and volume of 6 cubic inches? Round your answer to the tenths place. 0.5 inch 1.4 inches 3.0 inches 4.0 inches

OpenStudy (anonymous):

1.43239 in Given radius and volume calculate the height, slant height, lateral surface area and total surface area. Given r, V find h, s, L, A h = (3 * v) / (πr2)

OpenStudy (anonymous):

vol=1/3(pi)(r^2)(h)=6 (pi)r^2)(h)=18 3.14(4)h=18 12.56(h)=18 h=18/12.56=1.4 this is all said by helper99 i have laid it out differently in case it helps